# Modeling error correction with Lindblad dynamics and approximate channels

### v1

@@ Abstract
We analyze the performance of a quantum error correction code subject to physically-motivated noise modeled by a Lindblad master equation. Working within the code-capacity framework, we consider dissipative and coherent single-qubit terms and two-qubit crosstalk, studying how different approximations of the noise capture the performance of the five-qubit code. A composite-channel approximation where every noise term is considered separately, captures the behavior in many physical cases up to considerably-long timescales, and we analyze its eventual failure due to the effect of noncommuting terms. In contrast, we find that single-qubit approximations do not properly capture the error correction dynamics with two-qubit noise, even for short times. A Pauli approximation going beyond a single-qubit channel, is sensitive to the details of the noise, state, and decoder, and succeeds in many cases at short timescales relative to the noise strength, beyond which it fails. We calculate the code pseudo-threshold emerging within this model, and demonstrate how knowledge of the qubit parameters and connectivity can be used to design better decoders. These results shed light on the performance of error correction codes in the presence of realistic noise and can advance the ongoing efforts toward useful quantum error correction.
@@ I Introduction
Quantum computers represent a new paradigm in computing, holding great potential for applications in fields of science, technology, and industry. However, practical, large-scale quantum computers would require overcoming open challenges related to quantum error correction [1, 2, 3, 4, 5, 6, 7]. Error correction involves encoding the information contained in qubits into logical states belonging to a subspace of a higher-dimensional space, harnessing redundancy to facilitate error detection and the restoration of the original state. The states in this subspace are called logical states. With stabilizer error correction [8], logical states are identifiable by the measurement outcomes of a predefined set of observables. Those are referred to as stabilizers since they do not alter the logical states. A result of the stabilizers measurement is called a syndrome and when it is different from the one associated with a logical state, an error is identified. The syndrome serves as input to a decoder, which identifies the recovery process that will transform the corrupted state back into a logical state.
@@
When the recovery process successfully restores the original state, this is a success, and when it results in a different logical state, it is a failure. When analyzing the success rate, a common simplifying assumption is to model every single operation element of the noise as a Pauli operator (possibly multiqubit, though often single-qubit Pauli noise is assumed). The analysis proceeds by constructing a parity matrix, which conveys information about how the measurement outcomes of stabilizers are affected by specific Pauli errors. For any instance of Pauli noise (meaning a specification of the Pauli errors that occurred), the syndrome is obtained by multiplication of the parity matrix with a vector representing the error. This makes the first step, that of obtaining the syndrome, an efficient process that does not require simulating the logical state itself. As a result, this approach has been used extensively to study the behavior of error correction codes under various scenarios. These include, in particular, biased noise models [9, 10, 11, 12], correlated noise [13], and explorations of advantages of codes adapted (or tailored) to particular noise types [14, 15].
@@
However, Pauli noise channels are just a subset of completely positive and trace preserving (CPTP) quantum operations (or Kraus channels). Moreover, multiqubit noise channels can rarely be derived explicitly in physically-relevant cases [16]. In contrast, the generators of continuous dynamics of individual single qubits (1Q) and interacting two-qubit (2Q) pairs are often simpler to handle. Using the generators, the evolution of open systems is described by a master equation, and when the environment is memoryless (the dynamics being Markovian), the Lindblad master equation is the standard choice since it guarantees CPTP dynamics by construction [17, 18]. Lindbladian systems are extensively studied, with analytic, numerical and approximate solution methods continuously being improved [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30].
Despite the vast literature on open quantum systems, only a few existing works have compared error correction success rates between physically-motivated noise descriptions and their Pauli approximations. The majority of previous work considered only 1Q noise terms [31, 32, 33, 34]. Other works have focused on the effect of coherent errors (often considering analytically solvable cases), as in [35, 36, 37]. An earlier work devoted to linking a basic Lindblad model of a system of qubits to error correction success rates includes Ref. [38].
@@
In this paper, we study the accuracy of different approximations in modelling the behavior of a quantum error correction code when considering hardware-relevant Lindblad noise. We consider dynamics generated by 1Q coherent (unitary) and incoherent (dissipative) noise and also 2Q coherent interactions. We focus on the success rate of the five-qubit code [39], as it represents the smallest quantum error-correcting code capable of correcting any single 1Q error, and so is suitable as a toy-model code [37, 40]. We study the question of quantum-memory decoherence, with a logical code state evolving subject to the noise for varying durations. We compare the performance of the code with the state evolved under several approximations: (i) continuous Lindblad dynamics, the closest approach to realistic dynamics among those considered here (and the most computationally demanding one), (ii) an intermediate approach consisting of a composition of partial 2Q and 1Q Kraus channels, (iii) a Pauli approximation of the channels, and (iv) an attempt at an approximation using 1Q channels only.
@@
We perform our study within the framework of a code-capacity noise [41], which considers the logical state initialization, the syndrome measurements and the error recovery processes as ideal. This framework is often used (together with Pauli models) to calculate the error correction threshold of notable code families [42]. Although the current work is focused on a numerical study and the model relies on a simplified setup, we consider realistic parameters taken from an experimental characterization of state-of-the-art IBM Quantum superconducting-qubit devices accessible via the cloud [43], using the open-source packages Qiskit [44] and Qiskit Experiments [45]. We note also that the 1Q and 2Q noise terms that we study are relevant in other existing quantum hardware, rendering our conclusions applicable to various physical systems, as discussed in more detail later.
@@
The main results of this paper are as follows. The composite-channel approximation, which does not take into account the non-commutativity between noise terms, is a good approximation for relevant ranges of noise parameters, and up to considerably-long time scales. We point out the regime of its failure due to the neglect of noncommuting error terms. On the other hand, the validity of the Pauli approximation strongly depends on the noise parameters, initial state and decoder. For many cases it holds pretty well for short time scales, and diverges significantly for longer times. Therefore, before using this approximation it would seem necessary to check its validity for the specific details of the studied setup. We were unsuccessful at finding a 1Q approximation that holds even for short time scales. 1Q approximations differ significantly from the original dynamics, and so despite the benefit from the simplification of the simulation using this approximation, its discrepancy from the original dynamics could make it useless. Lastly, we study how knowledge of the details of the physical noise in the system, in particular the connectivity of the physical qubits, can be leveraged to improve the error correction code, by modifying the decoder to match the dominant noise.
@@ II The Lindbladian noise model
As the fundamental model for the qubits we consider the master equation describing the continuous-time evolution of a system’s density matrix $\rho(t)$ after being subject to the noisy channel for time $t$,
@@
where $\sigma^{a}_{i}$ is the Pauli matrix acting on qubit $i$ with $a\in\{x,y,z\}$, $I_{i}$ is the single-qubit identity matrix, the sum $\sum_{i}$ runs over all the qubits, $\sum_{\left\langle ij\right\rangle}$ runs over all coupled pairs, and $h_{i}$, $\zeta_{i,j}$ are coefficients in units of angular frequency. $H_{i}$ corresponds to a discrepancy between the actual qubit frequency and our estimated value of it (determined by the precision of the calibration experiment and ensuing drifts), and $V^{zz}_{ij}$ is a residual crosstalk between coupled qubits, resulting from uncontrolled interactions of ZZ type (also referred to as Ising coupling). We consider this coupling as a representative term of physical dynamics between idle connected qubits.
@@ III.1 Estimating error correction success rates
As discussed in Sec. I, when the noise is modeled as a Pauli noise, there is an efficient way to analyze the performance of the code, without the need of simulating a logical state. In contrast, when we are interested in a more complex physically-motivated noise described using continuous-time generators, there seems to be no way to analyze the code performance in a manner that is more efficient than simulating the logical state, which is how we proceed 111 A representation of the full channel acting on all qubits, as was done in [63], could be even harder to achieve numerically in the general case..
A brute-force simulation process would unfold as follows: It begins by selecting an initial logical state and initializing the qubits into this state. Subsequently, the noise is applied as dictated by the Lindbladian for a time duration $t$ corresponding to the time-scale of the pertinent noisy operator, resulting in the mixed state $\rho(t)$. Following this, for each potential syndrome outcome, the state is collapsed according to that particular syndrome, the relevant recovery operation is applied, and the resulting state is multiplied by the probability associated with that syndrome. Lastly, these states for all possible syndromes are summed. This process yields the mixed state of the system at the conclusion of the entire error correction procedure. The final step is evaluating the success rate of this procedure, achieved by computing the fidelity of the resulting state with the initial logical state. While the brute-force method is practical for small error correction codes, it becomes intractable for larger codes, primarily due to the exponential scaling of all possible syndromes and their corresponding recoveries.
In contrast, we present an alternative and more efficient approach to compute the code performance, addressing the issue of summing over all syndromes, though it does not eliminate the challenge of simulating the evolution of the full logical state. This method closely resembles the one employed in [31]. Every error correction code has a set of correctable errors that the code can rectify, $C\equiv\left\{C_{j}\right\}$. These errors can be chosen in such a way that their application to a logical code state results in orthogonal states, as discussed in [39]. Starting with an initial ideal logical state $|\psi_{0}\rangle$, we define $|\varphi_{j}\rangle=C_{j}|\psi_{0}\rangle$. If following the noise operation (before the syndrome extraction and recovery process) the system is in one of the states $|\varphi_{j}\rangle$ then the error correction procedure will successfully recover the initial state. If the system is in any other state, and since we are in the code-capacity model, the result will be a different logical state, which will be orthogonal to the initial logical state. Therefore, for a given mixed state $\rho(t)$, the probability of success is given by
$P_{C}=\sum_{\left\{\varphi_{j}\right\}}\left\langle\varphi_{j}|\rho(t)|\varphi_{j}\right\rangle=\sum_{C}\bra{\psi_{0}}C_{j}^{\dagger}\rho(t)C_{j}\ket{\psi_{0}}.$ (9)
Eq. (9) allows computing the success rate without simulating the syndrome measurement and recovery process, but rather only simulate the noisy evolution of the initial logical state. We define
@@
which is the failure probability of the recovery process of the original logical state. This quantity will serve as the figure of merit we calculate throughout this paper.
It is worth highlighting the distinctions between this method and the conventional approach for assessing the success rate of a code presented in Sec. I, in which any case where a logical operation is applied is categorized as a failure. In contrast, the approach that we take depends heavily on the initial state; If the error correction procedure results in a logical operation that does not alter the initial state, for instance, a logical $Z_{L}$ applied to the state $|0\rangle_{L}$, it is considered a success. In this sense, the conventional approach is a worst-case approach deeming every logical operator as a logical error, and this may lead to slight discrepancies between the two methods when applied on the same Pauli noise.
@@ III.2 Estimating a pseudo-threshold
In this framework, where the initial states are simulated explicitly, we choose to examine the pseudo-threshold on the average of the states
The averaging is done by computing $\eta$ separately for each state, and then taking the average of all states. Next, we compute the infidelity for the corresponding physical states, and average over these. Lastly, we find the crossing point of the logical $\eta$ and the physical infidelity.
@@
Nevertheless, for brevity we refer to the “restricted” code-capacity pseudo-threshold that we calculate simply as the pseudo-threshold, and note that it is informative for two reasons. The first is since it gives a rough idea of the order of magnitude of the parameter regime where this error correction is advantageous – thus roughly indicating the regime where the validity of the approximations is important for error correction. The second stems from the fact that the pseudo-threshold itself depends on the noise model, and so the difference between the pseudo-thresholds for the different approximations is another way to assess the validity of an approximation.

### v2

@@ Abstract
We analyze the performance of a quantum error correction code subject to physically motivated noise modeled by a Lindblad master equation. We consider dissipative and coherent single-qubit terms and two-qubit crosstalk, studying how different approximations of the noise capture the success rate of a code. Focusing on the five-qubit code and adapting it to partially correct two-qubit errors in relevant parameter regimes according to the noise model, we find that a composite-channel approximation where every noise term is considered separately captures the behavior in many physical cases up to long timescales, eventually failing due to the effect of noncommuting terms. In contrast, we find that single-qubit approximations do not properly capture the error correction dynamics with two-qubit noise, even for short times. A Pauli approximation going beyond a single-qubit channel is sensitive to the details of the noise, state, and decoder, and succeeds at short timescales relative to the noise strength, beyond which it fails. Furthermore, we point out a mechanism for a Pauli model failure where it underestimates the success rate of a code even with frequent syndrome projection and correction cycles. These results shed light on the performance of error correction in the presence of realistic noise and can advance the ongoing efforts towards useful quantum error correction.
@@ I Introduction
Quantum computers represent a new paradigm in computing, holding great potential for applications in fields of science, technology, and industry. However, practical, large-scale quantum computers would require overcoming open challenges related to quantum error correction [1, 2, 3, 4, 5, 6, 7, 8]. Error correction involves encoding the information into logical states belonging to a subspace of a higher-dimensional space, harnessing redundancy to facilitate error detection and the restoration of the original state. With stabilizer error correction [9], logical states are characterized by the measurement outcomes of a predefined set of observables, referred to as stabilizers since they do not alter the logical states. A result of the stabilizers measurement is called a syndrome and when it is different from the one associated with a logical state, an error is identified. The syndrome serves as input to a decoder, which identifies the recovery process that will transform the corrupted state back into a logical state.
@@
When the recovery process restores the original state, this is a success, and when it results in a different logical state, it is a failure. When analyzing the success rate, a common simplifying assumption is to model every single operation element of the noise as a Pauli operator (possibly multiqubit, though often single-qubit Pauli noise is assumed). The analysis proceeds by constructing a parity matrix, which conveys information about how the measurement outcomes of stabilizers are affected by specific Pauli errors. For any instance of Pauli noise (meaning a specification of the Pauli errors that occurred), the syndrome is obtained by multiplication of the parity matrix with a vector representing the error. This makes the first step, that of obtaining the syndrome, an efficient process that does not require simulating the logical state itself. As a result, this approach has been used extensively to study the behavior of error correction codes under various scenarios. These include, in particular, biased noise models [10, 11, 12, 13], correlated noise [14], and explorations of advantages of codes adapted (or tailored) to particular noise types [15, 16].
@@
However, Pauli noise channels are just a subset of completely positive and trace preserving (CPTP) quantum operations (or Kraus channels). Moreover, multiqubit noise channels can rarely be derived explicitly in physically-relevant cases [17]. In contrast, the generators of continuous dynamics of individual single qubits (1Q) and interacting two-qubit (2Q) pairs are often simpler to handle. Using the generators, the evolution of open systems is described by a master equation, and when the environment is memoryless (the dynamics being Markovian), the Lindblad master equation is the standard choice since it guarantees CPTP dynamics by construction [18, 19]. Lindbladian systems are extensively studied, with analytic, numerical and approximate solution methods continuously being improved [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32]. Nevertheless, only a few existing works have compared error correction success rates between physically-motivated noise descriptions and their Pauli approximations. The majority of previous work considered only 1Q noise terms [33, 34, 35, 36, 37]. Other works have focused on the effect of coherent errors (often considering analytically solvable cases), as in [38, 39, 40, 41], and crosstalk and gate errors have been simulated using Kraus channels with small codes [42, 5], or twirled to obtain Pauli channels as in [43]. Some works devoted to linking a basic Lindblad model of a system of qubits to error correction success rates include [44, 45] and some comparison to experiment in [46].
@@
In this paper, we study the accuracy of different approximations in modelling the behavior of a quantum error correction code when considering hardware-relevant Lindblad noise. We consider dynamics generated by 1Q coherent (unitary) and incoherent (dissipative) noise and also 2Q coherent interactions. We focus on the success rate of the five-qubit code [47], a toy-model code capable of correcting both bit- and phase-flip errors [40, 48], which we adapt to decode and correct 2Q errors (at the expense of some of the 1Q errors), according to the noise model. We study the question of quantum memory decoherence, with a logical code state evolving subject to the noise for varying durations. We compare the performance of the code with the state evolved under several approximations: (i) continuous Lindblad dynamics, the closest approach to realistic dynamics among those considered here (and the most computationally demanding one), (ii) an intermediate approach consisting of a composition of partial 2Q and 1Q Kraus channels, (iii) a Pauli approximation of the channels, and (iv) an attempt at an approximation using 1Q channels only.
@@
We perform our study within the framework of a code-capacity noise [49], which considers the logical state initialization, the syndrome measurements and the error recovery processes as ideal. This framework is often used (together with Pauli models) to calculate the error correction threshold of notable code families [50]. We also follow the dynamics of syndrome correction cycles, where we observe a nontrivial evolution even within the code-capacity framework. This simplified model allows investigating those dynamics and the validity of some common noise-approximation assumptions, which can be hard to isolate in more complex setups such as in the circuit-level model. For a realistic example, we take parameter values from an experimental characterization of currently deployed IBM Quantum superconducting-qubit devices accessible via the cloud [51], using the open-source packages Qiskit [52] and Qiskit Experiments [53]. The 1Q and 2Q noise terms that we study are relevant in other types of quantum hardware [43], as discussed also later.
@@
The main results of this paper are as follows. The composite-channel approximation, which is based on noise generators but does not take into account the non-commutativity between noise terms, is a good approximation for relevant ranges of noise parameters, and up to considerably long time scales. We point out the regime of its failure due to the neglect of noncommuting error terms. On the other hand, the validity of the multi-qubit Pauli approximation strongly depends on the noise parameters, initial state and decoder. For many cases it holds pretty well for short time scales, which are most relevant for error correction, and diverges significantly for longer times, and our study exemplifies a number of caveats of the Pauli approximation, whose relevance is broad and extends beyond the code capacity model. In particular, the Pauli noise model could fail in capturing effects of coherent noise and accumulation of phases in logical errors, even when frequently repeated syndrome cycles are applied. Therefore, before using this approximation it would seem necessary to check its validity for the specific details of the studied setup.
@@ II The Lindbladian noise model
As the fundamental model for the qubits’ dynamics we consider the master equation describing the continuous-time evolution of a system’s density matrix $\rho(t)$ when being subject to the noisy channel for time $t$,
@@
where $\sigma^{a}_{i}$ is the Pauli matrix acting on qubit $i$ with $a\in\{x,y,z\}$, $I_{i}$ is the single-qubit identity matrix, the sum $\sum_{i}$ runs over all the qubits, $\sum_{\left\langle ij\right\rangle}$ runs over all coupled pairs, and $h_{i}$, $\zeta_{i,j}$ are coefficients in units of angular frequency. $H_{i}$ corresponds to a discrepancy between the actual qubit frequency and our estimated value of it (determined by the precision of the calibration experiment and ensuing drifts), and $V^{zz}_{ij}$ is a residual crosstalk between coupled qubits, resulting from uncontrolled interactions of ZZ type (also referred to as Ising coupling). In physical systems the crosstalk would often couple nearest-neighbor qubits that would typically be data and parity (auxiliary) qubits, and parity qubits are lacking in our model to begin with. Nevertheless, effective longer-range interaction is also possible, and we consider this coupling as a representative term of coherent dynamics between qubits that are coupled either directly or indirectly. Some 2Q coherent dynamics is also expected when taking into account the over/under rotations of realistic 2Q gates.
@@ III.1 Estimating error correction success rates
As discussed in Sec. I, when the noise is modeled as a Pauli noise, there is an efficient way to analyze the performance of the code, without the need of simulating a logical state. In contrast, when we are interested in a more complex physically-motivated noise described using continuous-time generators, there seems to be no way to analyze the code performance in a manner that is more efficient than simulating the logical state, which is how we proceed 111 A representation of the full channel acting on all qubits, as was done in [71], could be even harder to achieve numerically in the general case.. The simulation begins by selecting an initial logical state and initializing the qubits into this state. We focus on the eigenstates of the logical Pauli operators, i.e., the states
Subsequently, the noise is applied as dictated by the Lindbladian for a time duration $t$ corresponding to the time scale of the pertinent noisy operator, resulting in the mixed state $\rho(t)$.
The next question is how to extract a useful metric from $\rho(t)$. As discussed further below, an explicit analysis that accounts for all possible syndromes and their corresponding recoveries could become challenging with large codes due to the exponential scaling of their number. Instead, we present in this subsection a simple and efficient approach to compute the code performance, suitable within the code capacity model and scalable to larger codes.
Every error correction code has a set of correctable errors that the code can rectify, $C\equiv\left\{C_{j}\right\}$. These errors can be chosen in such a way that their application to a logical code state results in orthogonal states, as discussed in [47]. Starting with an initial ideal logical state $|\psi_{0}\rangle$, we define $|\varphi_{j}\rangle=C_{j}|\psi_{0}\rangle$. If following the noise operation (before the syndrome extraction and recovery process) the system is in one of the states $|\varphi_{j}\rangle$ then the error correction procedure will successfully recover the initial state. If the system is in any other state, and since we are in the code capacity model, the result will be a different logical state, which will be orthogonal to the initial logical state.
Therefore, for a given state $\rho$, the probability of success is given by
$p_{C}=\sum_{\left\{\varphi_{j}\right\}}\left\langle\varphi_{j}|\rho|\varphi_{j}\right\rangle=\sum_{C}\bra{\psi_{0}}C_{j}^{\dagger}\rho C_{j}\ket{\psi_{0}}.$ (10)
Eq. (10) allows computing the success rate without simulating the syndrome measurement and recovery process, but rather only simulate the noisy evolution of the initial logical state. A similar formula has been used in [33]. We define
@@
which is the failure probability of the recovery process of the original logical state. We will use $\eta$ as the figure of merit we calculate and present in the analysis throughout most of this paper, except when stated otherwise, as discussed in more detail later.
It is worth highlighting the distinction between this method and the conventional approach for assessing the success rate of a code (based on a stochastic Pauli model) presented in Sec. I, in which every time a logical operation is applied is categorized as a failure. The approach that we take here depends on the initial state; if the error correction procedure results in a logical operation that does not alter the initial state, for example, a logical $Z_{L}$ applied to the state $|0\rangle_{L}$, it is considered a success. In this sense, the conventional approach is a worst-case approach deeming every logical operator as a logical error, and this may lead to slight discrepancies between the two methods when applied on the same Pauli noise.
@@ III.2 Estimating a pseudo-threshold
In this framework, where the initial states are simulated explicitly, we choose to examine the pseudo-threshold on the average of the states in Eq. (9). The averaging is done by computing $\eta$ separately for each state, and then taking the average of all states. Next, we compute the infidelity for the corresponding physical states, and average over these. Lastly, we find the crossing point of the logical $\eta$ and the physical infidelity.
@@
Nevertheless, for brevity we refer to the “restricted” code-capacity pseudo-threshold that we calculate simply as the pseudo-threshold, and note that it is informative for two reasons. The first reason is that it gives a rough idea of the order of magnitude of the parameter regime where this error correction is advantageous – thus roughly indicating the regime where the validity of the approximations is important for error correction. The second stems from the fact that the pseudo-threshold itself depends on the noise model, and so the difference between the pseudo-thresholds for the different approximations is another way to assess the validity of an approximation.
III.3 The logical density matrix
In this subsection we describe how to extract from the noisy state $\rho(t)$ more information than provided by $\eta$ defined in Eq. (11). Given a corrupt state $\rho$ and assuming ideal syndrome readout, for each potential syndrome outcome the state would be collapsed according to that particular syndrome and the relevant recovery operation applied. To construct the corresponding density matrix the resulting state is multiplied by the probability associated with that syndrome and the states for all possible syndromes are summed. This process yields the system state at the conclusion of the correction procedure,
$\rho\to\rho^{P}=\sum_{s}p_{s}\mathcal{C}_{s}(1/p_{s})P_{s}\rho P_{s}\mathcal{C}_{s}=\sum_{s}\mathcal{C}_{s}P_{s}\rho P_{s}\mathcal{C}_{s},$ (12)
with
$p_{s}={\rm tr}\{\rho P_{s}\},\quad P_{s}=\prod_{i}\left[\left(1\kern-2.5pt\text{l}+(-1)^{s_{i}}S_{i}\right)/2\right],$ (13)
where $S_{i}$ are the stabilizers of the code, and $\mathcal{C}_{s}$ is the correction operator corresponding to syndrome $s$ (composed of the bits $s_{i}$, to which corresponds the projector $P_{s}$).
For completeness, we note that the stabilizers of the five-qubit code that we investigate are
$S_{1}$ $=X_{0}Z_{1}Z_{2}X_{3}I_{4},$ (14)
$S_{2}$ $=I_{0}X_{1}Z_{2}Z_{3}X_{4},$ (15)
$S_{3}$ $=X_{0}I_{1}X_{2}Z_{3}Z_{4},$ (16)
$S_{4}$ $=Z_{0}X_{1}I_{2}X_{3}Z_{4},$ (17)
where $\{X,Y,Z\}$ denote the three Pauli matrices, and the indices indicate the qubit. The Pauli correction operators corresponding to each of the 16 possible syndromes can be found in Tab. 4 in App. L. The logical Pauli operators are
$Z_{L}=Z_{0}Z_{1}Z_{2}Z_{3}Z_{4},\quad X_{L}=X_{0}X_{1}X_{2}X_{3}X_{4}.$ (18)
The final step is to evaluate the success rate of this procedure, achieved by computing the fidelity of the resulting state with the ideal logical state |$\ket{\psi_{0}}$. Moreover, we can calculate also the matrix element of the state between |$\ket{\psi_{0}}$ and its orthogonal complement within the logical space, |$\ket{\psi_{0}^{\perp}}$, and define
[…]
To summarize, $\alpha$ is the infidelity of the corrected state with the ideal logical state, or in other words the probability of failure of the correction, and $\beta$ is a measure of coherent logical error that accumulates within the code space. Note that $\eta$ which was studied extensively above, is an approximation of $\alpha$ without explicitly performing the corrections in the circuit. Within the code-capacity model, when the projection onto the syndrome subspaces and the correction are both ideal, $\rho^{P}$ is guaranteed to be within the code space.